# Identifying heterogeneous micromechanical properties of biological tissues via physics-informed neural networks

### v1

@@ Abstract
The heterogeneous micromechanical properties of biological tissues have profound implications across diverse medical and engineering domains. However, identifying the full-field heterogeneous elastic properties of soft materials using traditional computational and engineering approaches is fundamentally challenging due to difficulties in estimating local stress fields. Recently, there has been a growing interest in using data-driven models to learn full-field mechanical responses such as displacement and strain from experimental or synthetic data. However, research studies on inferring the full-field elastic properties of materials, a more challenging problem, are scarce, particularly for large deformation, hyperelastic materials. Here, we propose a novel approach to identify the elastic modulus distribution in nonlinear, large deformation hyperelastic materials utilizing physics-informed neural networks (PINNs). We evaluate the prediction accuracies and computational efficiency of PINNs, informed by mechanic features and principles, across three synthetic materials with structural complexity that closely resemble real tissue patterns, such as brain tissue and tricuspid valve tissue. Our improved PINN architecture accurately estimates the full-field elastic properties, with relative errors of less than 5% across all examples. This research has significant potential for advancing our understanding of micromechanical behaviors in biological materials, impacting future innovations in engineering and medicine.
@@
Complex Materials, Micromechanical Properties, Physics-Informed Neural Networks, Heterogeneous Elastic Moduli
@@ 1 Introduction
Biological tissues exhibit a vast range of structural and mechanical heterogeneity [1, 2]. These properties have significant medical implications for both normal physiological function and disease. Research studies have shown that tissue biomechanical properties have a strong correlation to various cardiovascular diseases, including aortic disscection [3], aortic aneurysm [4, 5, 6], myocardial infarction [7], and degenerated heart valves [8]. Further, tissue mechanical properties also play a critical role in cancer invasion [9], tissue regenerative capacity [10, 11, 12, 13], host reaction upon implantation [14, 15, 16, 17], and fibrosis [18, 19]. Computational analysis, such as finite element modeling, has advanced our understanding of the physiological processes underlying these medical phenomena. However, despite the significant progress made in this field, many existing models rely on simplified constitutive models due to insufficient methodology to capture the complexities of material properties in soft tissues [20, 21]. This, in turn, impedes clinical translation of patient-specific computational analysis to inform medical decision making [22, 23, 24, 25].
The unique combination of soft material microstructure and properties not only provides indicators for diseases but also introduces various beneficial mechanical properties and functionalities that surpass those found in conventional materials. These desirable soft material characteristics have initiated researchers in additive manufacturing, materials science, and mechanics communities to explore the possibilities of biomimetic materials [26, 27], leading to innovative research in the design of artificial organs [28, 29, 30], soft robots [31, 32], and wearable devices [33]. Despite the apparent advantages of heterogeneous material design, including lower manufacturing costs, enhanced product usability, and improved end-user experience, there is a scarcity of investigations into the relationship between heterogeneous material properties and consequent functions, hindering the full realization of the functional potential of biomimetic materials [34].
@@
Atomic force microscopy (AFM) indentation and tensile test are two popular mechanical testing methods for determining the elastic properties of biological tissues. AFM is particularly effective for examining heterogeneous structures and properties of specimens at the cellular level [35, 36]. This method scans the tissue samples and applies perpendicular pressure to measure the resultant deformation at each probe location. Although this approach generates high spatial resolution maps of elastic moduli, the indentation loading conditions are not representative of physiological conditions of living tissues [37, 38], thus lacking relevance for the study of human disease and the mechanical behaviors of biological tissues.
On the other hand, in tensile testing, traction or deformation are applied to the boundaries of tissue samples. Digital image correlation (DIC) is utilized to capture the material deformation. Subsequent inverse finite element analysis and optimization methods are applied to identify the macroscopic constitutive parameters of the material [39, 40, 41]. While it is relatively straightforward to approximate stress distribution on the boundary of a testing specimen, extracting internal stress is non-trivial and often relies on a complex multiscale modeling approach [42, 43, 44]. Consequently, tensile testing has primarily been confined to the macroscopic homogeneous elastic properties of the material.
@@
Recently, researchers have begun to explore the application of data-driven methods to overcome the limitations mentioned above. In the context of mechanics, researchers have experimented with machine learning frameworks and architectures (e.g., convolutional neural networks, neural operators, and generative adversarial networks) to predict the mechanical responses and identify the elastic property map of heterogeneous materials of synthetical data [45, 46, 47, 48]. These methods have achieved high prediction accuracy. However, the drawback of data-driven approaches is that they typically require a large number of datasets, on the order of thousands to millions of image pairs (i.e., an input image representing data to be analyzed and an output image corresponding to the expected outcome), in order for the machine learning model to accurately learn their latent relation. Collecting a sufficient volume of images for successfully implementing a machine-learning model may require an impractical amount of effort and resources.
@@
Physics-informed machine learning [49, 50] provides an excellent alternative to traditional data-driven machine learning approaches in situations where the data is scarce [51, 52, 53]. Physics-informed neural networks (PINNs) have been utilized to determine the material parameters of homogenous structures within linear elastic, linear dynamic, hyperelastic, nonlinear-elastoplastic, and anisotropic constitutive models [54, 55, 56]. This approach has also been extended to identify the material parameters of structures with additional internal inclusions [57, 58, 59]. Recently, this has also been applied to linear elastic heterogeneous materials [60, 61]. It has been observed that PINNs struggle to converge in the presence of high-frequency functions or highly heterogeneous features [62, 63, 64]. As such, researchers have explored the application of Fourier-feature PINNs to simulate seismic wavefields [65], 1D radiation transport equations in heterogeneous media [66], and identify parameters in system-biology-informed ordinary differential equations (ODEs) problems [67, 68].
@@
The present research aims to tackle the more complex inverse problem that involves realistic tissue structural patterns and nonlinear material constitutive models. In particular, we aim to investigate the applicability and generality of PINNs for characterizing the nonlinear, hyperelastic properties of biological tissues such as brain tissue and tricuspid valve tissue, with an emphasis on discovering a high-fidelity map of complex elasticity fields in soft tissues using standard and Fourier-feature PINNs, as well as investigating the relationship between neural network architectures and prediction accuracy. This foundational work will be extensible to accurately estimate in vivo tissue properties from medical images and study the evolution of tissue microscopic elastic properties in response to mechanical loadings. We summarize the major contributions and significance of the current work as follows.
@@
  • We introduce a novel approach, using state-of-the-art PINN techniques, to discover the elastic modulus distribution in complex heterogeneous hyperelastic materials that have traditionally been challenging and mathematically ill-posed.
@@
  • We demonstrated that PINNs have remarkable potential for providing highly accurate parameter estimations using just one data sample. This is a significant breakthrough, as building an accurate machine-learning model through data-driven techniques can be time-consuming and impractical.
  • We analyzed the estimation accuracy of 20 neural network architectures, informed by mechanic features and principles, on three computational materials with complex local features. We identified a top-performing neural network architecture that consistently produces accurate estimations of the full-field elastic modulus. Further, the consistent accuracy of our selected top-performing neural network architecture suggests the generality of our method.
@@
  • This work provides a promising approach to studying the effect of material heterogeneity in their macroscopic behaviors and the evolution of material micromechanical properties in response to mechanical load. This is highly applicable to studying tissue growth and remodeling [69, 70, 71], detecting early mechanical markers of cancer cell invasion [72], and developing theoretical insights into the yield behavior of metallic alloys [73, 74].
@@ 2.1 Problem setup
Fig. 1 illustrates the structural pattern of three heterogeneous tissue specimens considered in the present work. The structural pattern of the first example was constructed using a Gaussian random field (GRF) [78] (Fig. 1A). The second example represents the cell bodies within a brain tissue [79] (Fig. 1B). The third example demonstrates the fibrosis network of the tricuspid valve leaflets [80] (Fig. 1C). We adopted a complexity measure based on image information theory, namely Delentropy [81, 82], to evaluate the complexity of tissue microstructural patterns. Delentropy captures the relative relation between the local and global features by analyzing the gradient vector field of the image. This measure provides a useful metric to evaluate the performance of machine learning models in the context of data complexity.
To generate reference strain data, we first assumed an elastic moduli distribution ($E$) for the tissue microstructure in the image. We then used the elastic distribution to compute the full-field strain data ($\epsilon_{xx}$, $\epsilon_{yy}$ and $\epsilon_{xy}$) using finite element analysis (FEA). Lastly, we integrated the full-field strain data in PINNs to guide the discovery of the heterogeneous elastic moduli distribution of the tissue microstructure, cell bodies, and fibrosis network in PINNs.
Figure 1: Heterogeneous material examples. The examples are presented in the undeformed configuration of unit square specimens. These examples represent the unknown elastic properties distribution of soft tissue that we seek to identify using PINNs. The structural complexity of the sample was quantified using Delentropy; higher values indicate greater complexity. (A) The structural pattern is synthetically generated using a GRF [78]. (B) The representation of the brain tissue microstructure was adapted from Koos et al. [79]. (C) The fibrosis network of the tricuspid valve was adapted from Weinberg et al. [80]. Equibiaxial deformation of magnitude $d$ was applied to the undeformed configuration using FEA to obtain the ground truth strain distribution. We prescribed 40% equibiaxial stretch to the specimen (i.e., 20% to each side of the boundary) and estimated the resulting Green-Lagrangian strain using FEniCS [83]. The strain data were used as ground truth reference data to identify the “unknown” elastic moduli in PINNs inversely.
@@ 2.2 PINN overview
Fig. 2 provides an overview of the network architectures in the present work. We examined the performance of four variants of PINNs, depicted in Figs. 2A to D. Figs. 2A and B represent standard PINNs architectures with and without enforcing boundary constraints, whereas Figs. 2C and D display Fourier-feature PINNs architectures with and without enforcing boundary constraints. Additionally, we identified five unique fully connected NN (FCNN) architectures (Fig. 2E) to investigate the prediction accuracy of the constitutive elastic properties within each PINNs variant. All simulations were performed on NVIDIA A100-SXM4-40GB GPUs. Detailed information related to Fourier-feature PINNs is described in Section 2.2.2.
@@ 2.2.1 PINN architectures
In the following, we briefly summarize the five FCNN architectures (Fig. 2E). We denote the output variables of FCNN I and FCNN II by $N_{u^{*}_{i}}$, $N_{P^{*}_{ij}}$, and $N_{E^{*}}$. Here, the output $N_{P^{*}_{ij}}$ is used to regulate the constitutive relation loss between the stress tensor $P^{*}_{ij}(u^{*}_{i},E^{*})$, computed from the estimated $u^{*}_{i}$ and $E^{*}$, and stress tensor $N_{P^{*}_{ij}}$ that directly estimated from the network. On the other hand, FCNN III, IV, and V are designed only to estimate $N_{u^{*}_{i}}$ and $N_{E^{*}}$ without ensuring a balanced constitutive relation loss.
@@
Regarding FCNN III, IV, and V, in FCNN III, we constructed a two-hidden layer shared network to connect the displacement components and elastic moduli. The shared network is subsequently split into three independent networks with three hidden layers with 25 neurons per layer to individual features in the displacement component and elastic moduli. FCNN IV consists of three independent networks, each with five hidden layers and 25 neurons per layer. FCNN V represents the simplest form of neural network architecture, in which one shared network is formulated to estimate $N_{u^{*}_{i}}$ and $N_{E^{*}}$.
In our analyses, we considered the strong form of linear momentum equations as the governing equation of the physical system
$P_{ij,i}=0,$
where $P$ is the first Piola-Kirchhoff stress described in Section. 5.1.
@@ 2.2.2 Fourier-feature PINNs
2.3 Overall findings on the computational experiments
@@
This section discusses the prediction accuracies of the proposed architectures in estimating the complex elastic property distributions of soft materials that undergo large deformation. The means and standard deviations of the $L^{2}$ relative errors of the estimated elastic modulus distribution of each example are shown in Figs. 3A to C. Each analysis was repeated with varying random seeds five times to ensure solution consistency. The FCNN that produces less than 5% average estimation errors are highlighted in bold. Furthermore, the architectures that consistently met the accuracy requirements across all three examples are underlined. The average computation time for each network architecture across the three examples and the five repeated tests are reported in Fig. 3D.
@@
Figure 3: Overall results. We conducted five evaluations for each PINN and FCNN architecture variant. (AC) The mean and standard deviations of $L^{2}$ relative errors of the PINN-estimated full-field elastic moduli. The network architectures that produced less than 5% $L^{2}$ relative errors are highlighted in bold. The architectures that consistently met the accuracy requirements across all three examples are underlined. Among the three network architectures that demonstrated high estimation accuracy across the examples, Network architecture IIB was the most computationally efficient. (D) The computational time for each PINN architecture.
@@
In the first example (Fig. 3A), PINN models A and B produced similar prediction accuracy across all five FCNN architectures. The difference in $L^{2}$ relative error between PINN models A and B was less than 1% for each FCNN architecture. On the contrary, four out of five FCNN architectures had higher $L^{2}$ relative errors in Fourier-feature PINN model C than D, suggesting, in this example, Fourier-feature PINNs provide more accurate estimates of the elastic parameter distribution when the boundary conditions were unconstrained.
In the second example (Fig. 3B), while FCNN III and V with PINN model A marginally outperformed those with PINN model B, the $L^{2}$ relative errors in FCNN architectures I, II, and IV were significantly higher with PINN models A than B, with differences in errors range from 1.35% to 5.99%. In contrast, Fourier-feature PINN model C produced comparable or better prediction than Fourier-feature PINN model D, with as much as 20.28% lower in $L^{2}$ relative error across the five FCNN architectures. In addition, FCNN III, IV, and V yielded the lowest $L^{2}$ relative errors (2.99%, 2.77%, and 2.61%, respectively) in Fourier-feature PINNs model C compared to the other models.
@@
In the third example (Fig. 3C), we observed similar $L^{2}$ relative error trends as those in Fig. 3B. Fourier-feature PINNs performed better than standard PINNs in this example. Moreover, Fourier-feature model C provided highly accurate estimations of the elastic parameter distribution with $L^{2}$ errors less than 5% across all FCNN architectures. In general, neural network architectures that enforce constitutive relations (FCNN I and II) demonstrate superior prediction accuracy in both standard PINNs and Fourier-feature PINNs when compared to architectures that do not enforce constitutive relations (FCNN III, IV, and V). This observation highlights the importance of considering the underlying physics and problem characteristics when designing and training PINNs.
@@
It is worth noting that Fourier-feature PINNs resulted in better estimations of elastic moduli in the second and third examples, but poorer estimations in the first example. This implies that the second and third examples display trigonometric features in their underlying microstructure, which are visibly apparent in the second and more subtle in the third. Therefore, the Fourier-feature transform provided additional guidance to the training process, subsequently improving the prediction accuracy. However, the first example did not seem to contain the same features. As a result, the Fourier-feature transform had a negative impact on the estimation of elastic property distribution.
@@
Out of the 20 network architectures tested, the network architectures IIB, IIC, and IVC were found to produce highly accurate estimations consistently across the three examples. The results suggest that partitioning output variables into separate independent networks, based on their shared features, may substantially enhance prediction accuracy. Among these three selected models, network architecture IIB provided accurate estimates of the elastic moduli, with average $L^{2}$ relative errors of 0.45%, 3.12%, and 3.05% for the three examples, while also being the fastest in terms of computational time.
@@
Regarding computational time, Fourier-feature PINNs took longer training time than standard PINNs. This is because of the additional neural network parameters introduced from the Fourier-feature mapping. Additionally, enforcing hard boundary conditions led to longer computational time in both standard and Fourier-feature PINNs. Among the PINN and Fourier-feature PINN models, model B required the least amount of training time overall, and among different architectures, FCNN architecture V required the least training time.
2.4 Displacement, strain, and elastic modulus estimation
Fig. 4 presents the estimated displacements, strain, and elastic modulus distribution of the three examples using network architecture IIB. The absolute relative errors between the PINN estimation and FEA ground truth are also provided for comparison. Despite the displacement boundaries being unconstrained in this network architecture, this PINN model demonstrated highly accurate estimation of the displacement, strain, and elastic parameter distribution across the three examples with relative errors up to $\mathcal{O}(10^{-2})$ for full-field displacements and strains and $\mathcal{O}(10^{-1})$ for the elastic moduli. These results indicate the consistency and generality of the proposed neural network architecture for estimating the mechanical properties of complex microstructures.
Figure 4: PINN predictions and pointwise error. The FEA ground truth, PINN estimation (denoted with superscript ***), and the absolute errors of displacement, strain, and elastic modulus distribution for examples (A), (B), and (C) are shown. The PINN architecture effectively resolves the local features of complex strain and elastic modulus fields. The absolute errors were up to $\mathcal{O}(10^{-2})$ for the displacement and up to $\mathcal{O}(10^{-1})$ for the strain and elasticity fields.
2.5 Convergence behavior of the PINN training

### v3

@@ Abstract
The heterogeneous micromechanical properties of biological tissues have profound implications across diverse medical and engineering domains. However, identifying full-field heterogeneous elastic properties of soft materials using traditional engineering approaches is fundamentally challenging due to difficulties in estimating local stress fields. Recently, there has been a growing interest in using data-driven models to learn full-field mechanical responses such as displacement and strain from experimental or synthetic data. However, research studies on inferring full-field elastic properties of materials, a more challenging problem, are scarce, particularly for large deformation, hyperelastic materials. Here, we propose a physics-informed machine learning approach to identify the elasticity map in nonlinear, large deformation hyperelastic materials. We evaluate the prediction accuracies and computational efficiency of physics-informed neural networks (PINNs) by inferring the heterogeneous elasticity maps across three materials with structural complexity that closely resemble real tissue patterns, such as brain tissue and tricuspid valve tissue. We further applied our improved architecture to three additional examples of breast cancer tissue and extended our analysis to three hyperelastic constitutive models: Neo-Hookean, Mooney Rivlin, and Gent. Our selected network architecture consistently produced highly accurate estimations of heterogeneous elasticity maps, even when there was up to 10% noise present in the training data.
@@
Physics-Informed Neural Networks, Complex Materials, Heterogeneous Mechanical Properties, Soft Tissues
@@ 1 Introduction
Biological tissues exhibit a vast range of structural and mechanical heterogeneity [1, 2]. These properties have significant medical implications for both normal physiological function and the treatment of diseases. Research studies have shown that tissue biomechanical properties are strongly correlated to cardiovascular diseases, including aortic disscection [3], aortic aneurysm [4, 5, 6], myocardial infarction [7], and degenerated heart valves [8]. Further, tissue and cell stiffness also play a critical role in identifying cancer invasion [9], tissue regenerative capacity [10, 11, 12, 13], host reaction upon implantation [14, 15, 16, 17], and fibrosis [18, 19]. Computational analysis, such as finite element modeling, provides a valuable non-invasive method to analyze the formation and progression underlying these medical phenomena [20, 21, 22, 23]. Yet, translating finite element computational analysis to inform medical decision-making still requires further model validation, which necessitates precise knowledge of the heterogeneous mechanical property distribution in biological tissues.
@@
There exist several methods for identifying the mechanical properties of materials, including atomic force microscopy (AFM) indentation, inverse finite element analysis, or virtual fields method. AFM is particularly effective for examining heterogeneous structures and properties of specimens at the cellular level [24, 25]. In particular, a cantilever probe scans the tissue samples and applies perpendicular pressure to measure the resulting deformation. The ratio of the applied pressure and deformation is subsequently used to approximate the tissue stiffness at each probed location. Although AFM produces high-resolution maps of elastic moduli, the indentation loading mode does not fully replicate the physiological conditions of living tissues [26, 27]. Therefore, the mechanical behaviors of biological tissues may be misrepresented.
In inverse finite element analysis, a finite element model is constructed to represent the geometry of the specimen being tested. The model is then subjected to appropriate boundary and loading conditions to simulate those in the experimental setup of a mechanical test. An initial guess of material constitutive parameters is applied to the finite element model and adjusted iteratively until the resulting mechanical responses agree with those measured in experiments. The estimation of mechanical parameters requires values of stress-strain pairs. While it is relatively straightforward to approximate stress distribution on the boundary of a testing specimen, extracting internal stress is non-trivial and often relies on a complex multiscale modeling approach [28, 29, 30]. As such, inverse finite element analysis has primarily been confined to deducing the macroscopic homogeneous elastic properties of the material [31, 32, 33]. The virtual fields method is another popular inverse method for identifying the full-field mechanical parameters of heterogeneous materials [34, 35]. The virtual fields method reconstructs the full-field stress data leveraging full-field displacement measurement provided by digital image correlation in conjunction with a preselected virtual displacement/velocity field function. The full-field material parameters are subsequently optimized by satisfying the Principle of Virtual Work equilibrium. It is important to note that the accuracy of parameter predictions is highly dependent on the selection of the virtual field function [36].
@@
Recently, there has been increased interest in applying data-driven methods to discover unknown material model parameters. Researchers have demonstrated high prediction accuracy of the mechanical responses and the elastic properties of materials derived from convolutional neural networks or neural operators  [37, 38, 39, 40]. However, the drawback of data-driven approaches is that they typically require a large number of datasets, on the order of thousands to millions of image pairs (i.e., an input image representing data to be analyzed and an output image corresponding to the expected outcome), in order for the machine learning model to learn their latent relation accurately. To this end, physics-informed machine learning [41, 42] has emerged as an excellent alternative to traditional data-driven machine learning approaches to reduce the high demands on training data while maintaining high prediction accuracy [43, 44, 45].
@@
Physics-informed neural networks (PINNs) have been utilized to determine the material parameters of homogenous structures within linear elastic, hyperelastic, and nonlinear-elastoplastic constitutive models [46, 47, 48, 49]. Following the initial success of PINNs, this approach has further been extended to identify the material parameters of structures with additional internal inclusions [50, 51, 52], as well as heterogeneous linear elastic materials [53, 54]. At times, PINNs may struggle to converge in the presence of high-frequency functions or highly heterogeneous features [55, 56, 57]. As such, the application of Fourier-feature PINNs in the context of system-biology-informed ordinary differential equations (ODEs) has also been explored [58, 59].
@@
The present research aims to tackle a complex inverse problem that involves realistic tissue structural patterns and nonlinear material constitutive models. In particular, we aim to investigate the applicability and generality of PINNs for characterizing the nonlinear, hyperelastic properties of biological tissues such as brain and tricuspid valve tissues, with an emphasis on discovering a high-fidelity map of complex elasticity fields in soft tissues, as well as investigating the relationship between neural network architectures and prediction accuracy. This foundational work will be extensible to accurately estimate in vivo tissue properties from medical images, where the full-field strain will be used as reference data, and study the evolution of tissue microscopic elastic properties in response to mechanical loadings. We summarize the major contributions and significance of the current work as follows.
@@
  • We introduce a physics-informed machine learning approach to discovering the elastic modulus distribution in complex heterogeneous hyperelastic materials that have traditionally been challenging and mathematically ill-posed.
  • Our improved neural network architecture consistently produces accurate estimates of full-field elasticity maps of complex material across multiple material constitutive models and biaxial stretch conditions.
@@
  • We demonstrated that PINNs have remarkable potential for providing highly accurate parameter estimations using just one data sample, even when there is a high level of noise in training data.
@@
  • This work provides a promising approach to studying the effect of material heterogeneity in their macroscopic behaviors and the evolution of material micromechanical properties in response to mechanical load. This is highly applicable to studying tissue growth and remodeling [60, 61, 62], and detecting early mechanical markers of cancer cell invasion [63].
@@ 2.1 Problem setup
Fig. 1 illustrates the structural pattern of three heterogeneous tissue specimens considered in the present work. The tissue specimens are subjected to equibiaxial stretch as it is one of the standard experimental techniques for studying the structural-mechanical relationship of soft tissues, including the brain, heart valve, and skin tissues [67, 68, 69]. The structural pattern of the first example was constructed using a Gaussian random field (GRF) [70] (Fig. 1A). The second example represents the cell bodies within a brain tissue [71] (Fig. 1B). The third example demonstrates the fibrous network of the tricuspid valve leaflets [72] (Fig. 1C). We adopted a complexity measure based on image information theory, namely Delentropy [73, 74], to evaluate the complexity of tissue microstructural patterns. Delentropy captures the relative relation between the local and global features by analyzing the gradient vector field of the image. This measure provides a useful metric to evaluate the performance of machine learning models in the context of data complexity.
To generate reference strain data, we first assumed an elastic moduli distribution ($E$) for the tissue microstructure in the image. We then applied equibiaxial stretch to the external boundary of the model and computed the full-field strain data ($\epsilon_{xx}$, $\epsilon_{yy}$, and $\epsilon_{xy}$) using finite element analysis (FEA). Lastly, we integrated the full-field strain data in PINNs to guide the discovery of the heterogeneous elastic moduli distribution of the tissue microstructure, cell bodies, and fibrous network in PINNs. Detailed description on the calculation of the elasticity map is discussed in Section 5.1. In the first part of the study (Section 2.3), we focused on identifying a suitable network architecture for identifying complex heterogeneous elasticity map. We applied 20% of outward prescribed displacements to the external boundary of the square specimens and assumed a plane strain, compressible Neo-Hookean material model. In the second part of our study (Section 2.4), we expanded our explorations to evaluate the performance of the selected PINN architecture under various material models, levels of equibiaxial stretch, and noise levels in the strain data.
Figure 1: Heterogeneous material examples. The examples are presented in the undeformed configuration of unit square specimens. These examples represent the unknown elastic properties distribution of soft tissue that we seek to identify using PINNs. The structural complexity of the sample was quantified using Delentropy; higher values indicate greater complexity. (A) The structural pattern is synthetically generated using a GRF [70]. (B) The representation of the brain tissue microstructure was adapted from Koos et al. [71]. (C) The fibrous network of the tricuspid valve was adapted from Weinberg et al. [72]. Equibiaxial deformation of magnitude $d$ was applied to the undeformed configuration using FEA to obtain the ground truth strain distribution. We prescribed 40% equibiaxial stretch to the specimen (i.e., 20% to each side of the boundary) and estimated the resulting Green-Lagrangian strain using FEniCS [75]. The strain data were used as ground truth reference data to identify the “unknown” elastic moduli in PINNs inversely.
@@ 2.2 PINN overview
In the current study, we evaluated the performance of four variations of PINNs, as shown in Figs. 2A to D. Figs. 2A and B represent standard PINNs architectures with and without boundary constraints, respectively. On the other hand, Figs. 2C and D display Fourier-feature PINNs architectures with and without boundary constraints, respectively. In addition, we examined the effect of five unique fully connected NN (FCNN) architectures (Fig. 2E) on the prediction accuracy of the constitutive elastic properties within each PINN variant. All simulations were performed on NVIDIA A100-SXM4-40GB GPUs. Detailed information related to Fourier-feature PINNs is described in Section 2.2.2.
@@ 2.2.1 PINN architectures
In the following, we briefly summarize the five FCNN architectures (Fig. 2E). The output variables of FCNN I and FCNN II are $N_{u^{*}_{i}}$, $N_{P^{*}_{ij}}$, and $N_{E^{*}}$. Here, the output $N_{P^{*}_{ij}}$ is used to govern the constitutive relation loss between the stress tensor $P^{*}_{ij}(u^{*}_{i},E^{*})$, computed from the estimated $u^{*}_{i}$ and $E^{*}$, and the direct stress tensor, $N_{P^{*}_{ij}}$, output from the network. On the other hand, FCNN III, IV, and V are designed only to estimate $N_{u^{*}_{i}}$ and $N_{E^{*}}$ without ensuring a balanced constitutive relation loss.
@@
Regarding FCNN III, IV, and V, in FCNN III, a two-hidden layer shared network is established to connect the displacement components and elastic moduli. The shared network is subsequently split into three independent networks with three hidden layers with 25 neurons per layer to individual features in the displacement component and elastic moduli. FCNN IV consists of three independent networks, each with five hidden layers and 25 neurons per layer. FCNN V represents the simplest form of neural network architecture, in which one shared network is formulated to estimate $N_{u^{*}_{i}}$ and $N_{E^{*}}$.
The strong form of linear momentum equations, $P_{ij,i}=0$, is considered as the governing equation of the physical system, wherein $P$ is the first Piola-Kirchhoff stress described in Section. 5.2.
@@ 2.2.2 Fourier-feature PINNs
2.3 Prediction accuracies of PINN architectures on elastic parameters estimation
@@
This section discusses the prediction accuracies of the proposed architectures in estimating the complex elastic property distributions of soft materials that undergo large deformation. The material constitutive model was assumed plane strain compressible Neo-Hookean with Poisson’s ratio $\nu=0.3$. The means and standard deviations of the $L^{2}$ relative errors of the estimated elastic modulus distribution of each example are shown in Figs. 3A to C. Each analysis was repeated five times with varying random seeds to ensure solution consistency. The FCNN that produced less than 5% average $L^{2}$ relative errors are highlighted in bold. Furthermore, the architectures that consistently met the accuracy requirements across all three examples are underlined. The average computation time for each network architecture across the three examples and the five repeated tests are reported in Fig. 3D.
@@
Figure 3: Effect of network architecture on PINN prediction. We conducted five evaluations for each PINN and FCNN architecture variant. (AC) The mean and standard deviations of $L^{2}$ relative errors of the PINN-estimated full-field elastic moduli. The network architectures that produced less than 5% $L^{2}$ relative errors are highlighted in bold. The architectures that consistently met the accuracy requirements across all three examples are underlined. Among the three network architectures that demonstrated high estimation accuracy across the examples, network architecture IIB was the most computationally efficient. (D) The computational time for each PINN architecture.
@@
In the first example (Fig. 3A), with standard PINNs, models A and B produced similar prediction accuracy across all five FCNN architectures. The difference in $L^{2}$ relative error between models A and B was less than 1% across all FCNN architectures. On the contrary, with Fourier-feature PINN, FCNN III, IV, and V had higher $L^{2}$ relative errors in model C than D. These results suggested that, in the GRF example, Fourier-feature PINNs provided more accurate estimates of the elastic parameter distribution when the boundary conditions were unconstrained and constitutive law was unenforced.
@@
Similar $L^{2}$ relative error patterns were observed in the second and third examples (Figs. 3B and  3C). We found that Fourier-feature PINNs outperformed standard PINNs in these two examples. Specifically, Fourier-feature model C yielded highly accurate estimates of the elastic parameter distribution with $L^{2}$ errors less than 5% across all FCNN architectures. Neural network architectures that enforce constitutive relations (FCNN I and II) demonstrated superior prediction accuracy in both standard PINNs and Fourier-feature PINNs when compared to those that do not enforce constitutive relations (FCNN III, IV, and V). This observation highlights the importance of considering the underlying physics and problem characteristics when designing and training PINNs.
@@
It is worth noting that Fourier-feature PINNs resulted in better estimates of the elasticity map in the second and third examples, but poorer estimates in the first example. This contradicting behavior suggests the existence of inherent trigonometric features in the full-field elasticity maps in the second and third examples, which were visibly apparent in the second and more subtle in the third. Therefore, the Fourier-feature transform provided additional guidance to the training process, subsequently improving the prediction accuracy. However, the first example did not seem to contain the same features. As a result, the Fourier-feature transform had a negative impact on the estimation of elastic property distribution.
@@
Out of the 20 network architectures tested, the network architectures IB, IIB, IC, and IVC were found to produce highly accurate estimations consistently across the three examples. The results suggest that partitioning output variables into separate independent networks, based on their shared features, may substantially enhance prediction accuracy. Among these selected models, network architecture IIB provided accurate estimates of the elastic moduli, with average $L^{2}$ relative errors of 0.41%, 3.45%, and 2.80% for the three examples respectively, while also being the fastest in terms of computational time.
@@
In terms of computational time, we observed that Fourier-PINNs require a longer duration of training in comparison to their standard PINN counterparts. This can be attributed to the introduction of additional neural network parameters through the Fourier-feature mapping. Additionally, enforcing hard boundary conditions led to longer computational time in both standard and Fourier-feature PINNs.
2.3.1 Comparison with the traditional inverse FEA method
Here, we provide a comparative study between PINNs and inverse FEA for the GRF example. In inverse FEA, we integrated a topology optimization method, the adjoint method [76], with FEA to minimize the $L^{2}$ relative errors between the ground truth and estimated strain tensor, subject to the constraint $1\leq E\leq 5$, the same as that in PINNs. Within this approach, the PDE and boundary condition losses were naturally optimized within each FEA iteration. The adjoint inverse analysis was performed using the Dolfin-adjoint software [77] with an Interior Point Optimizer [78]. The initial elasticity map was randomly generated. The maximum optimization iteration was set to 100, and the acceptable tolerance of the $L^{2}$ relative error for strains was set to $10^{-3}$. The results of our experiments are detailed in Appendix A. We found that the estimation accuracy using this optimization method is highly sensitive to the initial elasticity map.
The estimated strain tensor and full-field elasticity map using PINNs with network architecture IIB and the best elasticity map estimation from the adjoint method are reported in Fig. 4. Fig. 4A presents results from PINNs, while Fig. 4B demonstrates results using inverse FEA. As shown, the maximum absolute point-wise error of strains resulting from PINNs was two orders of magnitude lower than that from the adjoint method, and the full-field elasticity map was one order of magnitude lower. The $L^{2}$ relative error of the full-field elasticity map estimated by PINNs was 0.56%, while that estimated by the adjoint method was 29.22%. Note that, in the adjoint method, the optimization process terminated early due to solution divergence in the FEA.
Figure 4: Comparison of estimation accuracy using PINNs and the adjoint method. (A) In the results from PINNs, the maximum absolute point-wise errors of strains and the full-field elasticity map were 0.008 and 0.024. The $L^{2}$ relative error of the elasticity maps was 0.56%. (B) In the results from the adjoint method, the maximum absolute point-wise errors of strains and the full-field elasticity map were 0.25 and 0.741. The $L^{2}$ relative error of the elasticity maps was 29.22%. Our proposed PINN architecture provides significant improvement in predicting the GRF heterogeneous field compared to the adjoint method.
2.3.2 Convergence behavior of PINNs